OpenStudy (anonymous):

Solve the following equation for q. 5q + 2p + 4 = 3q – 8. a.q = -6 – 2p b. q = 2 – 2p c. q = 2 – p d. q = -6 – p

OpenStudy (anonymous):

can u medal me?

OpenStudy (anonymous):

at the right side please!!!

OpenStudy (anonymous):

can you answer my question first ? or nah ??

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

its hard wait

OpenStudy (anonymous):

uhmmm well if its wrong i wont medal you

OpenStudy (anonymous):

i think i know

OpenStudy (anonymous):

okay ?

OpenStudy (anonymous):

2(p + 3q + -4) + 5q(2p + -1q) = 0 Reorder the terms: 2(-4 + p + 3q) + 5q(2p + -1q) = 0 (-4 * 2 + p * 2 + 3q * 2) + 5q(2p + -1q) = 0 (-8 + 2p + 6q) + 5q(2p + -1q) = 0 -8 + 2p + 6q + (2p * 5q + -1q * 5q) = 0 -8 + 2p + 6q + (10pq + -5q2) = 0 Reorder the terms: -8 + 2p + 10pq + 6q + -5q2 = 0 Solving -8 + 2p + 10pq + 6q + -5q2 = 0 Solving for variable 'p'. Move all terms containing p to the left, all other terms to the right. Add '8' to each side of the equation. -8 + 2p + 10pq + 6q + 8 + -5q2 = 0 + 8 Reorder the terms: -8 + 8 + 2p + 10pq + 6q + -5q2 = 0 + 8 Combine like terms: -8 + 8 = 0 0 + 2p + 10pq + 6q + -5q2 = 0 + 8 2p + 10pq + 6q + -5q2 = 0 + 8 Combine like terms: 0 + 8 = 8 2p + 10pq + 6q + -5q2 = 8 Add '-6q' to each side of the equation. 2p + 10pq + 6q + -6q + -5q2 = 8 + -6q Combine like terms: 6q + -6q = 0 2p + 10pq + 0 + -5q2 = 8 + -6q 2p + 10pq + -5q2 = 8 + -6q Add '5q2' to each side of the equation. 2p + 10pq + -5q2 + 5q2 = 8 + -6q + 5q2 Combine like terms: -5q2 + 5q2 = 0 2p + 10pq + 0 = 8 + -6q + 5q2 2p + 10pq = 8 + -6q + 5q2 Reorder the terms: -8 + 2p + 10pq + 6q + -5q2 = 8 + -6q + 5q2 + -8 + 6q + -5q2 Reorder the terms: -8 + 2p + 10pq + 6q + -5q2 = 8 + -8 + -6q + 6q + 5q2 + -5q2 Combine like terms: 8 + -8 = 0 -8 + 2p + 10pq + 6q + -5q2 = 0 + -6q + 6q + 5q2 + -5q2 -8 + 2p + 10pq + 6q + -5q2 = -6q + 6q + 5q2 + -5q2 Combine like terms: -6q + 6q = 0 -8 + 2p + 10pq + 6q + -5q2 = 0 + 5q2 + -5q2 -8 + 2p + 10pq + 6q + -5q2 = 5q2 + -5q2 Combine like terms: 5q2 + -5q2 = 0 -8 + 2p + 10pq + 6q + -5q2 = 0

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

youre welcome